# FedHK-MVFC: Federated Heat Kernel Multi-View Clustering

### v1

@@ 3 Heat-Kernel Enhanced Multi-view Clustering
This section presents our proposed heat-kernel enhanced multi-view clustering framework. We first establish the mathematical foundations and notation used throughout our formulations, then develop our two novel clustering algorithms: Heat Kernel-Enhanced Multi-View Fuzzy Clustering (HK-MVFC) and its extension, Federated Heat Kernel Multi-View Fuzzy Clustering (FedHK-MVFC). Both algorithms are designed to effectively integrate information across heterogeneous data views in (non) distributed mechanisms while leveraging heat-kernel coefficients to enhance representational capacity.
@@ 3.1 Mathematical Preliminaries and Problem Formulation
Consider $X\in\mathbb{R}^{n\times D}$ as a multi-view dataset consisting of $s$ separate views, denoted by ${X^{h}}=\left\{{x_{1}^{h},\ldots,x_{n}^{h}}\right\}_{h=1}^{s}$. Each $x_{i}^{h}=\left\{{x_{ij}^{h}}\right\}_{j=1}^{{d_{h}}}$ signifies the $i$-th data point in the $h$-th view, and the total dimensions are such that $\sum\nolimits_{h=1}^{s}{{d_{h}}=D}$. The data matrix $X$ is divided into $c$ clusters over all $s$ views, represented by a global membership matrix ${\mu^{*}}={\left[{\mu_{ik}^{*}}\right]_{n\times c}}$, where each element $\mu_{ik}^{*}\in\left[{0,1}\right]$ adheres to the rule $\sum\nolimits_{k=1}^{c}{\mu_{ik}^{*}=1}$. The weights of the views are given by the vector $v={\left[{{v_{h}}}\right]_{1\times s}}$, with ${v_{h}}\in\left[{0,1}\right]$ and ensuring $\sum\nolimits_{h=1}^{s}{{v_{h}}=1}$. Cluster centroids in each view $h$ are indicated by ${A^{h}}={\left[{a_{kj}^{h}}\right]_{c\times{d_{h}}}}$.
@@ 3.2 Heat-Kernel Coefficients
Central to our approach is the formulation of heat-kernel coefficients (H-KC), which enables the transformation of Euclidean distance in the original feature space to an exponential kernel distance. For the $j$-th feature in the $h$-th view, we define H-KC $\delta_{ij}^{h}$ using two alternative estimators:
@@ 3.4 Heat Kernel-Enhanced Multi-View Fuzzy Clustering
To address the challenges of clustering heterogeneous multi-view data, we propose the Heat Kernel-Enhanced Multi-View Fuzzy Clustering (HK-MVFC) algorithm. This approach integrates heat-kernel coefficients into a fuzzy clustering framework, enabling more effective capture of complex relationships between data points and cluster centers across multiple views.
The foundation of our approach is the transformation of conventional Euclidean distance metrics into exponential kernel distances. To achieve this, we introduce a novel Kernel Euclidean Distance (KED) formulation:
@@
${J_{HK-MVFC}}\left({V,U,A}\right)=\sum\limits_{h=1}^{s}{v_{h}^{\alpha}}\sum\limits_{i=1}^{n}{\sum\limits_{k=1}^{c}{{{\left({\mu_{ik}^{*}}\right)}^{m}}}}d_{\exp\left({ik,j}\right)}^{h}$ (6)
@@
A critical aspect of our approach is the use of a unified membership matrix $U=\left[{\mu_{ik}^{*}}\right]$ across all views. Unlike approaches that compute separate memberships for each view and subsequently merge them, HK-MVFC enforces consistency across views from the outset, resulting in a more coherent clustering structure.
@@ 3.5 Optimization Framework: HK-MVFC
The objective function of proposed HK-MVFC in Eq. 6 cannot be minimized directly due to its non-convex nature. We therefore employ an alternating optimization approach, iteratively updating the membership matrix $U$, the view weights $V$, and the cluster centers $A$.
@@
${{\tilde{J}}_{HK-MVFC}}={J_{HK-MVFC}}\left({V,U,A}\right)+\sum\limits_{i=1}^{n}{{\lambda_{1i}}}\left({\sum\limits_{k=1}^{c}{\mu_{ik}^{*}-1}}\right)+{\lambda_{2}}\left({\sum\limits_{h=1}^{s}{{v_{h}}-1}}\right)$ (7)
@@
The necessary conditions for minimizing the objective function $J_{HK-MVFC}$ in Eq. 6 yield the following update rules:
@@
We start by formulating the update equation for the membership coefficients $\mu_{ik}^{*}$. This is done by computing the partial derivative of the Lagrangian presented in Eq. 7 with respect to $\mu_{ik}^{*}$:
$\frac{{\tilde{\partial}{J_{HK-MVFC}}}}{{\partial\mu_{ik}^{*}}}=m{\left({\mu_{ik}^{*}}\right)^{m-1}}\sum\limits_{h=1}^{s}{v_{h}^{\alpha}}d_{\exp\left({ik,j}\right)}^{h}+{\lambda_{1i}}=0$ (11)
$\Rightarrow\mu_{ik}^{*}={\left({-\frac{{{\lambda_{1i}}}}{m}}\right)^{{{\left({m-1}\right)}^{-1}}}}{\left({\sum\limits_{h=1}^{s}{v_{h}^{\alpha}d_{\exp\left({ik,j}\right)}^{h}}}\right)^{-{{\left({m-1}\right)}^{-1}}}}$ (12)
By imposing the condition $\sum\nolimits_{k=1}^{c}{\mu_{ik}^{*}=1}$, we obtain:
$\sum\limits_{k=1}^{c}{{{\left({-\frac{{{\lambda_{1i}}}}{m}}\right)}^{{{\left({m-1}\right)}^{-1}}}}}{\left({\sum\limits_{h=1}^{s}{v_{h}^{\alpha}d_{\exp\left({ik,j}\right)}^{h}}}\right)^{-{{\left({m-1}\right)}^{-1}}}}=1$ (13)
$\Rightarrow{\left({-\frac{{{\lambda_{1i}}}}{m}}\right)^{{{\left({m-1}\right)}^{-1}}}}=\frac{1}{{\sum\limits_{k^{\prime}=1}^{c}{{{\left({\sum\limits_{h=1}^{s}{v_{h}^{\alpha}d_{\exp\left({ik^{\prime},j}\right)}^{h}}}\right)}^{-{{\left({m-1}\right)}^{-1}}}}}}}$ (14)
Substituting Eq. 14 into Eq. 12 yields Eq. 8.
@@
To find the cluster centers $a_{kj}^{h}$, we differentiate the HK-MVFC objective in Eq. 6 with respect to $a_{kj}^{h}$:
$\frac{{\partial{J_{HK-MVFC}}}}{{\partial a_{kj}^{h}}}=v_{h}^{\alpha}\sum\limits_{i=1}^{n}{{{\left({\mu_{ik}^{*}}\right)}^{m}}}\frac{{\partial d_{\exp\left({ik,j}\right)}^{h}}}{{\partial a_{kj}^{h}}}$ (15)
$=v_{h}^{\alpha}\sum\limits_{i=1}^{n}{{{\left({\mu_{ik}^{*}}\right)}^{m}}}\frac{\partial}{{\partial a_{kj}^{h}}}\left\{{1-\exp\left({-\sum\limits_{j=1}^{{d_{h}}}{\delta_{ij}^{h}{{\left({x_{ij}^{h}-a_{kj}^{h}}\right)}^{2}}}}\right)}\right\}$ (16)
$=-v_{h}^{\alpha}\sum\limits_{i=1}^{n}{{{\left({\mu_{ik}^{*}}\right)}^{m}}}\frac{\partial}{{\partial a_{kj}^{h}}}\exp\left({-\sum\limits_{j=1}^{{d_{h}}}{\delta_{ij}^{h}{{\left({x_{ij}^{h}-a_{kj}^{h}}\right)}^{2}}}}\right)$ (17)
$=2\delta_{ij}^{h}v_{h}^{\alpha}\sum\limits_{i=1}^{n}{{{\left({\mu_{ik}^{*}}\right)}^{m}}\exp\left({-\sum\limits_{j=1}^{{d_{h}}}{\delta_{ij}^{h}{{\left({x_{ij}^{h}-a_{kj}^{h}}\right)}^{2}}}}\right)}\left({x_{ij}^{h}-a_{kj}^{h}}\right)$ (18)
Setting the derivative in Eq. 18 to zero and solving for $a_{kj}^{h}$ gives Eq. 9. Then, we compute the partial derivative of the Lagrangian in Eq. 7 with respect to $v_{h}$} to update the weight function for views $v_{h}$}.
$\frac{{\partial{{\tilde{J}}_{HK-MVFC}}}}{{\partial{v_{h}}}}=\alpha v_{h}^{\alpha-1}\sum\limits_{i=1}^{n}{\sum\limits_{k=1}^{c}{{{\left({\mu_{ik}^{*}}\right)}^{m}}}}d_{\exp\left({ik,j}\right)}^{h}+{\lambda_{2}}=0$ (19)
$\Rightarrow{v_{h}}={\left({-\frac{{{\lambda_{2}}}}{\alpha}}\right)^{{{\left({\alpha-1}\right)}^{-1}}}}{\left({\sum\limits_{i=1}^{n}{\sum\limits_{k=1}^{c}{{{\left({\mu_{ik}^{*}}\right)}^{m}}d_{\exp\left({ik,j}\right)}^{h}}}}\right)^{-{{\left({\alpha-1}\right)}^{-1}}}}$ (20)
From constraint $\sum\nolimits_{h=1}^{s}{{v_{h}}=1}$, we derive Eq. 20.

### v2

@@ 3 Heat-Kernel Enhanced Multi-view Clustering
This section presents our proposed heat-kernel enhanced multi-view clustering framework. We first establish the mathematical foundations and notation used throughout our formulations, then develop our two novel clustering algorithms: Heat Kernel-Enhanced Multi-View Fuzzy Clustering (HK-MVFC) and its extension, Federated Heat Kernel Multi-View Fuzzy Clustering (FedHK-MVFC). HK-MVFC algorithm is designed to effectively integrate information across heterogeneous data views in (non-distributed) mechanisms while leveraging heat-kernel coefficients to enhance representational capacity. In particular, we aim to address the following research problems, including how we can effectively integrate multiple heterogeneous data views into a unified clustering framework that captures complex relationships among data points across views and how we can leverage heat-kernel coefficients to enhance the representational capacity of clustering algorithms, particularly in capturing the intrinsic geometries of data manifolds.
@@ 3.1 Mathematical Preliminaries and Problem Formulation
We begin by establishing the mathematical framework and notation that will be used throughout our formulation. This formulation serves as the foundation for both the centralized HK-MVFC and federated FedHK-MVFC algorithms.
3.1.1 Multi-View Data Representation
Consider a multi-view dataset $X\in\mathbb{R}^{n\times D}$ consisting of $n$ data samples observed across $s$ distinct views. Each view $h\in\{1,2,\ldots,s\}$ captures a different perspective or modality of the underlying data, represented as:
$X^{h}=\{x_{1}^{h},x_{2}^{h},\ldots,x_{n}^{h}\}\quad\text{for }h=1,2,\ldots,s$ (1)
where each data point $x_{i}^{h}\in\mathbb{R}^{d_{h}}$ in view $h$ is characterized by $d_{h}$ features:
$x_{i}^{h}=[x_{i1}^{h},x_{i2}^{h},\ldots,x_{id_{h}}^{h}]^{T}\quad\text{for }i=1,2,\ldots,n$ (2)
The total dimensionality of the multi-view dataset is given by $D=\sum_{h=1}^{s}d_{h}$, representing the aggregated feature space across all views.
3.1.2 Clustering Parameters
The fundamental objective is to partition the $n$ data samples into $c$ clusters, where the clustering structure is consistent across all views. This clustering is characterized by several key mathematical objects:
Unified Membership Matrix: The clustering assignment is represented by a global fuzzy membership matrix $U^{*}\in\mathbb{R}^{n\times c}$, where each element $\mu_{ik}^{*}$ denotes the degree of membership of data point $i$ to cluster $k$:
$U^{*}=[\mu_{ik}^{*}]_{n\times c},\quad\text{where }\mu_{ik}^{*}\in[0,1]$ (3)
The membership values satisfy the probabilistic constraint that ensures each data point’s total membership across all clusters sums to unity:
$\sum_{k=1}^{c}\mu_{ik}^{*}=1\quad\text{for all }i=1,2,\ldots,n$ (4)
View Weight Vector: To account for the varying importance and informativeness of different views, we introduce an adaptive view weight vector $V\in\mathbb{R}^{1\times s}$:
$V=[v_{1},v_{2},\ldots,v_{s}],\quad\text{where }v_{h}\in[0,1]$ (5)
These weights are normalized to ensure a valid probability distribution over views:
$\sum_{h=1}^{s}v_{h}=1$ (6)
Higher values of $v_{h}$ indicate that view $h$ is more informative for the clustering task, enabling the algorithm to automatically emphasize more reliable data sources.
Cluster Center Matrix: For each view $h$, the cluster prototypes are represented by a cluster center matrix $A^{h}\in\mathbb{R}^{c\times d_{h}}$:
$A^{h}=[a_{kj}^{h}]_{c\times d_{h}}\quad\text{for }h=1,2,\ldots,s$ (7)
where $a_{kj}^{h}$ represents the $j$-th feature of the $k$-th cluster center in view $h$. The complete set of cluster centers across all views is denoted as $\mathcal{A}=\{A^{1},A^{2},\ldots,A^{s}\}$.
3.1.3 Problem Statement
Given the multi-view dataset $X=\{X^{1},X^{2},\ldots,X^{s}\}$ and a specified number of clusters $c$, our objective is to determine the optimal clustering parameters $(U^{*},V,\mathcal{A})$ that minimize a unified objective function. This objective function will integrate information across all views while accounting for the intrinsic geometric structure of the data through heat kernel-enhanced distance measures.
The mathematical framework established here provides the foundation for developing both centralized and federated multi-view clustering algorithms that can effectively handle heterogeneous data sources while preserving data privacy in distributed settings.
@@ 3.2 Heat-Kernel Coefficients
Central to our approach is the formulation of heat-kernel coefficients (H-KC), which enables the transformation of Euclidean distance in the original feature space into an exponential kernel distance. For the $j$-th feature in the $h$-th view, we define H-KC $\delta_{ij}^{h}$ using two alternative estimators:
@@ 3.4 Heat Kernel-Enhanced Multi-View Fuzzy Clustering
To address the challenges of clustering heterogeneous multi-view data, we propose the Heat Kernel-Enhanced Multi-View Fuzzy Clustering (HK-MVFC) algorithm. This approach integrates H-KC into a fuzzy clustering framework, enabling more effective capture of complex relationships between data points and cluster centers across multiple views.
The foundation of our approach is the transformation of CED metrics into exponential kernel distances. To achieve this, we introduce a novel Kernel Euclidean Distance (KED) formulation:
@@
${J_{HK-MVFC}}\left({V,U^{*},A}\right)=\sum\limits_{h=1}^{s}{v_{h}^{\alpha}}\sum\limits_{i=1}^{n}{\sum\limits_{k=1}^{c}{{{\left({\mu_{ik}^{*}}\right)}^{m}}}}d_{\exp\left({ik,j}\right)}^{h}$ (13)
@@
A critical aspect of our approach is the use of a unified membership matrix $U^{*}=\left[{\mu_{ik}^{*}}\right]$ across all views. Unlike approaches that compute separate memberships for each view and subsequently merge them, HK-MVFC enforces consistency across views from the outset, resulting in a more coherent clustering structure.
@@ 3.5 Optimization Framework: HK-MVFC
The objective function of proposed HK-MVFC in Eq. 13 cannot be minimized directly due to its non-convex nature. We therefore employ an alternating optimization approach, iteratively updating the membership matrix $U^{*}$, the view weights $V$, and the cluster centers $A$.
@@
${{\tilde{J}}_{HK-MVFC}}={J_{HK-MVFC}}\left({V,U^{*}A}\right)+\sum\limits_{i=1}^{n}{{\lambda_{1i}}}\left({\sum\limits_{k=1}^{c}{\mu_{ik}^{*}-1}}\right)+{\lambda_{2}}\left({\sum\limits_{h=1}^{s}{{v_{h}}-1}}\right)$ (14)
@@
The necessary conditions for minimizing the objective function $J_{HK-MVFC}$ in Eq. 13 yield the following update rules for the membership matrix $\mu^{*}$, cluster centers $a$, and view weights $v$:
@@
Membership Update: Eq. 15 updates the membership degree $\mu_{ik}^{*}$ of data point $x_{i}$ to cluster $k$, where the weighted distances across all $s$ views are aggregated using view weights $v_{h}^{\alpha}$, and the fuzzifier parameter $m$ controls the degree of membership overlap.
@@
Cluster Center Update: Eq. 16 computes the $j$-th feature of cluster center $a_{k}$ in view $h$ as a weighted average of data points, where weights combine membership degrees $(\mu_{ik}^{*})^{m}$, view importance $v_{h}^{\alpha}$, and exponential kernel terms $\exp(-\delta_{ij}^{h}\|x_{i}^{h}-a_{k}^{h}\|^{2})$.
@@
View Weight Update: Eq. 17 determines the importance weight $v_{h}$ for view $h$, inversely proportional to the total weighted distance in that view, where $\alpha$ controls the distribution of weights across views (normalized such that $\sum_{h=1}^{s}v_{h}=1$).
Remark 1.
These update rules form an iterative alternating optimization scheme: given current estimates of cluster centers and view weights, Eq. 15 updates memberships; given memberships and view weights, Eq. 16 updates centers; and given memberships and centers, Eq. 17 updates view weights. This process continues until convergence.
@@
We establish the necessary optimality conditions for the HK-MVFC objective function using Lagrangian optimization and derive the closed-form update rules for each parameter set.
Part I: Membership Matrix Update Rule Derivation
@@
To derive the update equation for the membership coefficients $\mu_{ik}^{*}$, we begin by computing the partial derivative of the Lagrangian in Eq. 14 with respect to $\mu_{ik}^{*}$. This derivative represents the marginal change in the objective function when we adjust the membership of data point $i$ to cluster $k$.
$\frac{\partial\tilde{J}_{HK-MVFC}}{\partial\mu_{ik}^{*}}$ $=m(\mu_{ik}^{*})^{m-1}\sum_{h=1}^{s}v_{h}^{\alpha}d_{\exp(ik,j)}^{h}+\lambda_{1i}=0$ (18)
where the first term $m(\mu_{ik}^{*})^{m-1}$ arises from the power rule applied to the fuzzified membership $(\mu_{ik}^{*})^{m}$, and the summation $\sum_{h=1}^{s}v_{h}^{\alpha}d_{\exp(ik,j)}^{h}$ represents the weighted contribution of all views to the distance between data point $i$ and cluster center $k$.
Solving Eq. 18 for $\mu_{ik}^{*}$ yields:
$\mu_{ik}^{*}$ $=\left(-\frac{\lambda_{1i}}{m}\right)^{(m-1)^{-1}}\left(\sum_{h=1}^{s}v_{h}^{\alpha}d_{\exp(ik,j)}^{h}\right)^{-(m-1)^{-1}}$ (19)
Here, the exponent $(m-1)^{-1}=\frac{1}{m-1}$ transforms the equation from its derivative form back to the original membership variable. The negative sign in the Lagrange multiplier term ensures that memberships decrease as distances increase, which is the desired clustering behavior.
To determine the Lagrange multiplier $\lambda_{1i}$, we impose the normalization constraint $\sum_{k=1}^{c}\mu_{ik}^{*}=1$, which ensures that the total membership of data point $i$ across all clusters sums to unity (a fundamental requirement in fuzzy clustering):
$\sum_{k=1}^{c}\left(-\frac{\lambda_{1i}}{m}\right)^{(m-1)^{-1}}\left(\sum_{h=1}^{s}v_{h}^{\alpha}d_{\exp(ik,j)}^{h}\right)^{-(m-1)^{-1}}$ $=1$ (20)
$\Rightarrow\left(-\frac{\lambda_{1i}}{m}\right)^{(m-1)^{-1}}$ $=\frac{1}{\sum_{k^{\prime}=1}^{c}\left(\sum_{h=1}^{s}v_{h}^{\alpha}d_{\exp(ik^{\prime},j)}^{h}\right)^{-(m-1)^{-1}}}$ (21)
The denominator in Eq. 21 serves as a normalization factor that ensures all memberships sum to one. Substituting this expression back into Eq. 19 yields the final membership update rule in Eq. 15, where each membership is inversely proportional to the weighted distance raised to the power $\frac{1}{m-1}$.
Part II: Cluster Centers Update Rule Derivation
@@
To find the optimal cluster centers $a_{kj}^{h}$, we differentiate the HK-MVFC objective function in Eq. 13 with respect to $a_{kj}^{h}$. This derivative measures how the objective function changes when we adjust the $j$-th feature of cluster center $k$ in view $h$:
$\frac{\partial J_{HK-MVFC}}{\partial a_{kj}^{h}}$ $=v_{h}^{\alpha}\sum_{i=1}^{n}(\mu_{ik}^{*})^{m}\frac{\partial d_{\exp(ik,j)}^{h}}{\partial a_{kj}^{h}}$ (22)
where $v_{h}^{\alpha}$ weights the contribution of view $h$, and $(\mu_{ik}^{*})^{m}$ weights each data point’s contribution by its membership to cluster $k$.
Expanding the kernel distance derivative:
$\frac{\partial d_{\exp(ik,j)}^{h}}{\partial a_{kj}^{h}}$ $=\frac{\partial}{\partial a_{kj}^{h}}\left\{1-\exp\left(-\sum_{j=1}^{d_{h}}\delta_{ij}^{h}(x_{ij}^{h}-a_{kj}^{h})^{2}\right)\right\}$ (23)
$=-\frac{\partial}{\partial a_{kj}^{h}}\exp\left(-\sum_{j=1}^{d_{h}}\delta_{ij}^{h}(x_{ij}^{h}-a_{kj}^{h})^{2}\right)$ (24)
The exponential term captures the heat kernel transformation, where $\delta_{ij}^{h}$ are the heat-kernel coefficients that adapt to local data geometry. Applying the chain rule:
$\frac{\partial J_{HK-MVFC}}{\partial a_{kj}^{h}}$ $=2\delta_{ij}^{h}v_{h}^{\alpha}\sum_{i=1}^{n}(\mu_{ik}^{*})^{m}\exp\left(-\sum_{j=1}^{d_{h}}\delta_{ij}^{h}(x_{ij}^{h}-a_{kj}^{h})^{2}\right)(x_{ij}^{h}-a_{kj}^{h})$ (25)
The factor of 2 arises from differentiating the squared term $(x_{ij}^{h}-a_{kj}^{h})^{2}$, while the exponential term weights contributions based on the heat kernel similarity between data points and cluster centers. The term $(x_{ij}^{h}-a_{kj}^{h})$ indicates the direction of adjustment: positive when data points are larger than the center, negative otherwise.
Setting Eq. 25 to zero and rearranging yields Eq. 16, which shows that cluster centers are weighted averages of data points, where weights combine fuzzy memberships $(\mu_{ik}^{*})^{m}$, view importance $v_{h}^{\alpha}$, and heat kernel similarities $\exp(-\delta_{ij}^{h}\|x_{i}^{h}-a_{k}^{h}\|^{2})$.
Part III: View Weights Update Rule Derivation
To update the view weights $v_{h}$, we compute the partial derivative of the Lagrangian in Eq. 14 with respect to $v_{h}$. This derivative quantifies how the objective function changes when we adjust the importance assigned to view $h$:
$\frac{\partial\tilde{J}_{HK-MVFC}}{\partial v_{h}}$ $=\alpha v_{h}^{\alpha-1}\sum_{i=1}^{n}\sum_{k=1}^{c}(\mu_{ik}^{*})^{m}d_{\exp(ik,j)}^{h}+\lambda_{2}=0$ (26)
The term $\alpha v_{h}^{\alpha-1}$ comes from differentiating $v_{h}^{\alpha}$ with respect to $v_{h}$, while the double summation $\sum_{i=1}^{n}\sum_{k=1}^{c}(\mu_{ik}^{*})^{m}d_{\exp(ik,j)}^{h}$ represents the total weighted distance contribution of view $h$ across all data points and clusters.
Solving for $v_{h}$:
$v_{h}$ $=\left(-\frac{\lambda_{2}}{\alpha}\right)^{(\alpha-1)^{-1}}\left(\sum_{i=1}^{n}\sum_{k=1}^{c}(\mu_{ik}^{*})^{m}d_{\exp(ik,j)}^{h}\right)^{-(\alpha-1)^{-1}}$ (27)
The exponent $(\alpha-1)^{-1}=\frac{1}{\alpha-1}$ transforms the derivative relationship back to the view weight variable. The negative exponent on the distance summation ensures that views with smaller total distances (better clustering quality) receive higher weights.
Applying the normalization constraint $\sum_{h=1}^{s}v_{h}=1$ determines the Lagrange multiplier $\lambda_{2}$ and yields the final view weight update in Eq. 17. This adaptive weighting mechanism automatically assigns higher importance to more informative views, where informativeness is measured by the view’s ability to create compact, well-separated clusters.
The alternating optimization of these three parameter sets (memberships, cluster centers, and view weights) constitutes the complete HK-MVFC algorithm, with each update rule derived from necessary optimality conditions of the constrained optimization problem. ∎